Shift report 20000729 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sat Jul 29 2000 - 16:43:00 EDT

  • Next message: http account: "Shift report 20000730 16:00-24:00"

    ------------------------------------------------------------
    Date: Sat Jul 29 16:42:59 2000, Shift: 08:00-16:00
    Supervisor: kh, with: tk, fv
    Log book#3 pages 107 - 109
    
    Shift Summary:
    ------------------------------------------------------------
    MCR worked on injecting and accelerating beam during
    most of the shift.  Finally about 3:30pm they achieved
    stored and cogged beam at gamma=70.
    We observe some collision candidates, but rate is very 
    low.  At shift end they are steering cogged beam to STAR.
    
    Started run 1972.  FV working on timing in FEH at shift
    end.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sat Jul 29 2000 - 16:46:26 EDT